Circular No. 7933

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002dx IN UGC 12861 B. Swift and W. D. Li, University of California at Berkeley (UCB),

report the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered KAIT CCD images taken on July 8.5 (mag ∼ 18.8) and 9.5 UT (mag ∼ 18.9). SN 2002dx is located at α = 23^{h}57^{m}06^{s}.73, δ = +29^{o}50′15′′.5 (equinox 2000.0), which is 4′′.0 east and 5′′.3 south of the nucleus of UGC 12861. A KAIT image taken on 2001 Nov. 19.2 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVA 2002dy IN MCG –01-59-24 Swift and Li also report the LOTOSS discovery, on unfiltered KAIT

images taken on July 8.5 and 9.5 UT, of an apparent supernova (mag ∼ 17.0) located at α = 23^{h}28^{m}23^{s}.27, δ = −2^{o}48′17′′.0 (equinox 2000.0), which is 21′′.4 west and 16′′.0 south of the nucleus of MCG –01-59-24. A KAIT image taken on 2001 Nov. 2.2 showed nothing at this position (limiting mag ∼ 19.5).

S. Jha, T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrogram (range 370–750 nm; low signal-to-noise ratio) of SN 2002dy, obtained by M. Calkins on July 10.47 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-II supernova. The spectrum, taken during poor weather conditions, shows a flat continuum with broad Hα and Hβ emission (FWHM 3100 km/s; reminiscent of type-IIn supernovae) at the redshift of the host galaxy (9904 km/s; NASA/IPAC Extragalactic Database). The Hβ profile shows a hint of an absorption trough, at an expansion velocity of 9200 km/s.

SUPERNOVAE 2002dv AND 2002dw A. V. Filippenko, R. Chornock, and R. J. Foley, UCB, report that in-

spection of CCD spectra (range 310–1000 nm), obtained on July 11 UT with the Shane 3-m reflector at Lick Observatory, reveals that both SN 2002dv (IAUC 7930) and SN 2002dw (IAUC 7932) are young type-II supernovae. SN 2002dv exhibits strong Hα emission with a weaker P-Cyg absorption component. Hβ and Fe II lines with P-Cyg profiles are also present. SN 2002dw has a nearly featureless continuum, possibly contaminated by hot stars in an underlying H II region, as well as weak, broad Hα emission.

Corrigenda. On IAUC 7932, line 6, for June 4.3; 19.4 on June read July 4.3; 19.4 on July

2002 July 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green